What is the hypotenuse of a right triangle whose catheters are 3 and 4? Imercian reader
1
Expert's answer
2020-04-06T18:17:36-0400

By the Pythagorean theorem we have

"a^2+b^2=c^2" (1)

in the right triangle "ABC" (see the picture below). Here is "a" and "b" are legs and "c" is the hypotenuse. From (1) it follows that

"c^2=3^2+4^2=25,"

"c=5."